# Three-term Machin-type formulae

### v4

@@ 1. Introduction
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ divides $x_{i}^{2}+1$ for $i=1,2,3$, then $m_{1}<m_{2}<1.342\times 10^{34},x_{i}<\exp(8380114833)$ and $\left\lvert y_{i}\right\rvert<9.152\times 10^{16}$.
@@
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ does not divide $x_{i}^{2}+1$ for some $i$, then $m_{1}<1.017\times 10^{24}$, $m_{2}<\exp(191323)$, $x_{i}<\exp(3.217\times 10^{12})$, and $\left\lvert y_{i}\right\rvert<8.334\times 10^{18}$. Moreover, $m_{2}<e^{40000}$ if additionally $m_{1}\geq 3.537\times 10^{22}$.
@@
    If $m_{1}m_{2}<10^{10}$, then $x_{i}<\exp(8813999998)$, and $\left\lvert y_{i}\right\rvert<4.508\times 10^{18}$.
@@
These upper bounds themselves do not give finiteness of $m_{1}$ and $m_{2}$. However, noting that $r\neq 0$, which gives $\left\lvert\sum_{i}y_{i}\arctan(1/x_{i})\right\rvert\geq\pi/4$, the first case can be easily settled using these upper bounds. In order to settle the second case, we additionally need an upper bound $m_{2}$ in terms of $m_{1}$. This can be done using a lower bound for a quantity of the form $y\arctan(1/x)-r\pi/2$, which gives a linear form of two logarithms.
Our PAGI-GP script is available from
@@ 3. A lower bound for linear forms of three logarithms
Now we set $\rho$ according to Table 1 and take $a_{1}=\rho\pi/2$ and $a_{3}=\rho\pi/2+\log m_{2}$. Moreover, we can take $a_{2}=2\rho\arctan(Y_{1}/X_{1})+\log m_{1}$ if $m_{1}\leq 41$ and $a_{2}=\rho\pi/2+\log m_{1}$. Hence, we have $\rho>4$, $a_{1}\geq 2\pi$, $a_{2}\geq 6\arctan(1/4)+\log 17=4.79304\cdots$. Moreover, we see that $a_{2}a_{3}\geq(9.062\arctan(1/4)+\log 37)(2.2655\pi+\log 270270277)=135.5131\cdots$. We note that in the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*}\})$, we have $a_{2}a_{3}\geq(\log(53)+\rho\pi/2)(\log(10^{10}/53)+\rho\pi/2)$ since $(\log x+\rho\pi/2)(\log(10^{10})-\log x+\rho\pi/2)$ is increasing in the range $1\leq x\leq 10^{5}$.
@@
(35) $c_{i}^{\prime}=(\chi_{1}C_{1})^{2/3}+\frac{C_{2}C_{0}^{1/3}C_{1}^{2/3}}{\log^{1/3}\rho}+\max\left\{\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}},\sqrt{\frac{C_{0}C_{1}}{a_{5-i}^{2}a_{1}\log\rho}}\right\}$
@@
We write $i=k+1,j=k+2$ for brevity and put $R_{i1}=\left\lfloor(\chi^{2}K^{2}a_{i+1}a_{i+2}/a_{i}^{2})^{1/3}\right\rfloor$ for $i=1,2,3$.
Since $\chi_{1}>0.6$, $C_{1}>6$, $K>540000$, $C_{0}\leq 3.5$, and $a_{1}a_{2}a_{3}>20$ by our settings and $b_{0}\geq B_{0}=e^{20}$ by assumption, we have $((1-1/K)\chi_{1}C_{1})^{2/3}>C_{0}^{1/2}/(2\sqrt{a_{1}a_{2}a_{3}\log^{1/2}b_{0}})$ and therefore
(37) $R_{k+1,1}+1>(\chi^{2}(K-1)^{2}a_{k}a_{k+2}/a_{k+1}^{2})^{1/3}>\sqrt{\frac{C_{0}a_{k+2}\log b_{0}}{4a_{k+1}}}$
and, similarly,
(38) $R_{k+2,1}+1>\sqrt{\frac{C_{0}a_{k+1}\log b_{0}}{4a_{k+2}}}.$
Hence, we obtain $4(R_{k+1,1}+1)(R_{k+2,1}+1)>L$ to confirm (ii).
@@
(40) $\begin{split}(R_{11}+1)(R_{21}+1)(R_{31}+1)>&\leavevmode\nobreak\ (\chi_{1}C_{1}a_{1}a_{2}a_{3})^{2}\log^{3}b_{0}>K\chi_{1}^{2}C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0}\\ >&\leavevmode\nobreak\ K\max\{R_{11}+R_{21}+1,R_{21}+R_{31}+1,R_{31}+R_{11}+1\}.\end{split}$
@@
In each case, noting that $a_{k+2}>a_{k+1}>a_{k}$, we have $R_{k,j}\geq R_{k+1,j}\geq R_{k+2,j}$ for $j=1,3$ and $R_{k+2,j}$ can be minorized by replacing $m_{2}$ by $\tilde{m}_{2}$ and then $m_{1}$ by $\tilde{m}_{1}$. In the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$, $R_{k+2,j}$ can be minorized by replacing $m_{2}$ by $10^{10}/m_{1}$ and then taking $m_{1}=53$. Moreover, we observe that $R_{i2}\geq\left\lfloor(2K^{2}a_{k}a_{i}/a_{j}^{2})^{1/3}\right\rfloor$, which can be minorized by the same way. Hence, we have $R_{1}+R_{2}+R_{3}\geq R_{0}$ for $i=1,2,3$ for constants $K_{0}$ and $L_{0}$ in Table 2 in each case.
@@
(49) $C_{3}=\frac{g_{0}C_{0}}{\log\rho}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}\log^{2}B_{0}}+C_{3}^{\prime}\right).$
Now we would like to show that $\log b\leq\log b_{0}$.
We see that $R_{i}-1=R_{i1}+R_{i2}+R_{i3}<c_{i}^{\prime}a_{i+1}a_{i+2}\log b_{0}$ for $i=1,2,3$. Hence, we obtain
@@
observing that
(54) $\frac{2+6\pi^{-2}+\log K}{3K(K-1)}-\frac{1+2\log(2\pi K/e^{3/2})}{K-1}<0$
for $K>1$. Hence, we have
@@
(60) $\left(1-\frac{1}{K}\right)\left(1-\frac{1}{L}-\frac{1}{3L^{2}}\right)C_{0}C_{1}\geq\frac{2\log(C_{0}(C_{1}a_{1}a_{2}a_{3})^{2}\log^{2}b_{0}/\log\rho)}{a_{1}a_{2}a_{3}\log^{2}b_{0}}+C_{1}+C_{3}$
@@
(78) $\begin{split}&\rho_{1}\left\lvert\log\alpha_{i}^{\prime}\right\rvert+2\operatorname*{h}(\alpha_{i}^{\prime})<u_{1}\left(\log m_{i}+\frac{\rho_{1}\pi}{2}\right)+\frac{u_{i+1}\rho_{1}\pi}{2}\\ &<(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}\left(\left(\log m_{i}+\frac{\rho_{1}\pi}{2}\right)\rho+\left(\log m_{i}+\frac{\rho\pi}{2}\right)\frac{\rho_{1}\pi}{2}\right)\\ &=(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}((\rho+\rho_{1})\log m_{i}+\rho\rho_{1}\pi)\pi/2.\end{split}$
@@
(93) $\begin{split}{b^{\prime}}=&\frac{b_{1}}{9.05\pi+2D\operatorname*{h}(\alpha)}+\frac{b_{2}}{9.05\pi},\\ D=&[{\mathbb{Q}}(\alpha):{\mathbb{Q}}]/2,\\ a=&9.05\pi+2D\operatorname*{h}(\alpha),\\ H=&\max\{1000,D,D(\log{b^{\prime}}+2.96)+0.01\}.\end{split}$
@@
(94) $\log\left\lvert\Lambda_{1}\right\rvert>-1.8793aH^{2}.$
@@
Then, we can confirm that $0.2790.2432<\sqrt{k}<$ 0.279, $H>7.5$, $15\leq L<0.92h$, and $K>700$ and then the condition (2) of Theorem 1 of [6]. Moreover, we can also confirm that $R_{2}-1<2a_{2}H^{2}<b_{1}$ and $S_{2}-1<2a_{1}H^{2}<b_{2}$. Hence the condition (1) of Theorem 1 of [6] also holds. Then, confirming that $\sqrt{k}L<0.19665H$, Theorem 1 of [6] yields that $\log\left\lvert\Lambda_{1}\max\{LSe^{LS\left\lvert\Lambda_{1}\right\rvert/(2b_{2})}/(2b_{2}),LRe^{LR\left\lvert\Lambda_{1}\right\rvert/(2b_{1})}/(2b_{1})\}\right\rvert>-1.8792aH^{2}$ and then $\log\left\lvert\Lambda_{1}\right\rvert>-1.8793aH^{2}$. For the detail of the proof, see [21]. ∎
@@ 5. Proof of the Theorem
If $m_{2}\geq 10^{34}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{34})$ again, we have $\left\lvert y_{i}\right\rvert<9.098\times 10^{16}$ and $\log x<8380114833$. If $10^{32}\leq m_{2}<10^{34}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{32})$, we have $\left\lvert y_{i}\right\rvert<9.152\times 10^{16}$ and $\log x<8373165769$. If $10^{24}\leq m_{2}<10^{32}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{24})$, we have $\left\lvert y_{i}\right\rvert<8.752\times 10^{16}$ and $\log x<7706422610$. If $m_{2}<10^{24}$, then, applying Theorem 4.1 again with $\tilde{m}_{1}=53$ and $\tilde{m}_{2}=m_{1}^{*}$, we have $\left\lvert y_{i}\right\rvert<7.799\times 10^{16}$ and $\log x<5456058110$. Hence, in any case, we have $y_{i}<9.152\times 10^{16}$ and $x_{i}<\exp(8380114833)$.
@@
We begin by showing that $m_{1}<M_{1}$ with $M_{1}=1.017\times 10^{24}$. Assume that $m_{2}\geq m_{1}\geq M_{1}$. Now, Theorem 4.1 allows us to take $A_{4}=1752.029$, $A_{5}=10545512.8\cdots$, and $\rho=7.292$.
@@
Now, if $m_{1}\leq 6.5\times 10^{8}G^{2}$, then, (121) gives $m_{1}<6.5\times 10^{8}A_{4}^{2}(\log^{2}m_{1})\log^{4}(10545513\log m_{1})$ and therefore we must have $m_{1}<1.012\times 10^{24}<M_{1}$. Thus, we may assume that $m_{2}>m_{1}>6.5\times 10^{8}G^{2}$.
@@
(125) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{m_{1}}}>(1-10^{-9})\left(\frac{\pi}{4}-7.9\times 10^{-5}\right)>0.78531.$
@@
(130) $\log\left\lvert\Lambda_{1}\right\rvert>-1.8793aH^{2},$
@@
(131) $\begin{split}a=&9.05\pi+\log m_{1}\\ H=&\max\left\{1000,2.97+\log\left(\frac{\left\lvert r\right\rvert}{9.05\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9.05\pi}\right)\right\}.\end{split}$
@@
Since $m_{1}\geq M_{1}$, we have $0.3774942.97+\log\psi_{3}(m_{1})\leq 2.97+\log\psi_{3}(M_{1})<-$ 0.377494. Hence, if $H>1000$, then
@@
We observe that the right hand side of this inequality is at least $\log(A_{4}\log M_{1})\log^{2}(10545513\log M_{1})>4000$ to obtain
@@
whether $H>1000$ or not. With the aid of these inequalities, we deduce from (129) and (130) that
@@
(140) $A_{4}(\alpha\log m_{2})\log^{2}(10545513\alpha\log m_{1})<0.78515\sqrt{m_{1}}$
@@
Now we show that $m_{2}<\exp(191323)$. Assume that $m_{2}\geq\exp(191323)$. If $m_{1}\geq M_{1}^{\prime}:=3.537\times 10^{22}$, then, since $m_{2}>e^{40000}$, we iterate the argument that we have just done with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ to obtain (138) and (139) with $28.85883$, $82.15586$, and $281.5761$ in place of $103.5805$, $268.0433$, and $1246.539$. Eventually, we obtain
(141) $A_{4}(\alpha\log m_{2})\log^{2}(6665783\alpha\log m_{1})<0.785395\sqrt{m_{1}}$
with $A_{4}=1675.249$ instead of (140). On the other hand, if $m_{1}\leq 5\times 10^{14}G^{2}$, then, (121) gives $G<A_{4}(\log^{2}(5\times 10^{14}G^{2}))\log^{4}(6665783\log(2.56\times 10^{10}G^{2}))$ and therefore $G<3\times 10^{12}$, which implies that $m_{2}>e^{40000}>5\times 10^{14}G^{2}$. Hence, in any case, we must have $m_{2}>5\times 10^{14}G^{2}$ to obtain (124) and then (127) with $10^{-7}$ in place of $7.9\times 10^{-5}$ and $0.785398$ in place of $0.78531$, contrary to (141). Hene, we must have $m_{1}<M_{1}^{\prime}$ when $m_{2}\geq e^{40000}$.
@@
Moreover, if $5\leq m_{1}\leq 41$ and $m_{2}\geq e^{40000}$, then we obtain (137) with $A_{4}$ taken from Table 4 with $\tilde{m}_{2}=e^{40000}$, which is incompatible with (127) when $m_{2}\geq e^{77847}$.
@@
Hence, we must have $m_{1}<1.017\times 10^{24}$ and $m_{2}<\exp(191323)$. Moreover, if $m_{1}\geq 3.537\times 10^{22}$, then $m_{2}<e^{40000}$. If $m_{1}\geq 10^{14}$ and $m_{2}\geq e^{40000}$, then, Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ yields that $\log x\leq(\log 2+k_{i}\log m_{1}+\ell_{i}\log m_{2})/2<3.217\times 10^{12}$, $x_{i}<\exp(3.217\times 10^{12})$, and $\left\lvert y_{i}\right\rvert\leq 2FG<8.334\times 10^{18}$. Similarly, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})$ according to Table 6 in each case, we can confirm that $x_{i}<\exp(3.217\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<8.334\times 10^{18}$ whenever $m_{1}\geq 53$.

### v5

@@ 1. Introduction
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ divides $x_{i}^{2}+1$ for $i=1,2,3$, then $m_{1}<m_{2}<1.342\times 10^{34},x_{i}<\exp(8380114833)$, and $\left\lvert y_{i}\right\rvert<9.152\times 10^{16}$.
@@
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ does not divide $x_{i}^{2}+1$ for some $i$, then $m_{1}<2.549\times 10^{24}$, $m_{2}<\exp(295490)$, $x_{i}<\exp(5.07\times 10^{12})$, and $\left\lvert y_{i}\right\rvert<1.316\times 10^{19}$. Moreover, $m_{2}<e^{40000}$ if additionally $m_{1}\geq 9.203\times 10^{22}$.
@@
    If $m_{1}m_{2}<10^{10}$, then $x_{i}<\exp(8813999998)$ and $\left\lvert y_{i}\right\rvert<4.508\times 10^{18}$.
@@
These upper bounds themselves do not give finiteness of $m_{1}$ and $m_{2}$. However, noting that $r\neq 0$, which gives $\left\lvert\sum_{i}y_{i}\arctan(1/x_{i})\right\rvert\geq\pi/4$, the first case can be easily settled using these upper bounds. In order to settle the second case, we additionally need an upper bound $m_{2}$ in terms of $m_{1}$. This can be done using a lower bound for a quantity of the form $y\arctan(1/x)-r\pi/2$, which gives a linear form of two logarithms. Computation was done by PAGI-GP. Out PARI-GP script is available from
@@ 3. A lower bound for linear forms of three logarithms
Now we set $\rho$ according to Table 1 and take $a_{1}=\rho\pi/2$ and $a_{3}=\rho\pi/2+\log m_{2}$. Moreover, we can take $a_{2}=2\rho\arctan(Y_{1}/X_{1})+\log m_{1}$ if $m_{1}\leq 41$ and $a_{2}=\rho\pi/2+\log m_{1}$ otherwise. Hence, we have $\rho>4$, $a_{1}\geq 2\pi$, $a_{2}\geq 6\arctan(1/4)+\log 17=4.79304\cdots$. Moreover, we see that $a_{2}a_{3}\geq(9.062\arctan(1/4)+\log 37)(2.2655\pi+\log 270270277)=135.5131\cdots$. We note that in the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$, we have $a_{2}a_{3}\geq(\log(53)+\rho\pi/2)(\log(10^{10}/53)+\rho\pi/2)$ since $(\log x+\rho\pi/2)(\log(10^{10})-\log x+\rho\pi/2)$ is increasing in the range $1\leq x\leq 10^{5}$.
@@
(35) $c_{i}^{\prime}=(\chi_{1}C_{1})^{2/3}+\frac{C_{2}C_{0}^{1/3}C_{1}^{2/3}}{\log^{1/3}\rho}+\max\left\{\frac{2^{1/3}C_{1}^{2/3}}{\log^{1/3}B_{0}},\sqrt{\frac{C_{0}C_{1}}{a_{1}\log\rho}}\right\}$
@@
We put $R_{i1}=\left\lfloor(\chi^{2}K^{2}a_{i+1}a_{i+2}/a_{i}^{2})^{1/3}\right\rfloor$ for $i=1,2,3$. Since $\chi_{1}>0.6$, $C_{1}>6$, $K>540000$, $C_{0}\leq 3.5$, and $a_{1}a_{2}a_{3}>20$ by our settings and $b_{0}\geq B_{0}=e^{20}$ by assumption, we have $((1-1/K)\chi_{1}C_{1})^{2/3}>C_{0}^{1/2}/(2\sqrt{a_{1}a_{2}a_{3}\log b_{0}})$ and therefore
(37) $R_{i,1}+1>(\chi^{2}K^{2}a_{i+1}a_{i+2}/a_{i}^{2})^{1/3}>\sqrt{\frac{C_{0}a_{1}a_{5-i}\log b_{0}}{4a_{i}}}$
for $i=1,2$. Hence, we obtain $4(R_{2,1}+1)(R_{3,1}+1)>L$ to confirm (ii).
@@
(39) $\begin{split}(R_{11}+1)(R_{21}+1)(R_{31}+1)>&\leavevmode\nobreak\ \chi_{1}^{2}K^{2}>K\chi_{1}^{2}C_{1}a_{1}a_{2}a_{3}\log^{2}b_{0}\\ >&\leavevmode\nobreak\ K\max\{R_{11}+R_{21}+1,R_{21}+R_{31}+1,R_{31}+R_{11}+1\}.\end{split}$
@@
In each case, noting that $a_{k+2}>a_{k+1}>a_{k}$, we have $R_{k,j}\geq R_{k+1,j}\geq R_{k+2,j}$ for $j=1,3$ and $R_{k+2,j}$ can be minorized by replacing $m_{2}$ by $\tilde{m}_{2}$ and then $m_{1}$ by $\tilde{m}_{1}$. In the case $(\tilde{m}_{1},\tilde{m}_{2})=(53,m_{1}^{*})$, $R_{k+2,j}$ can be minorized by replacing $m_{2}$ by $10^{10}/m_{1}$ and then taking $m_{1}=53$. Moreover, we observe that $R_{i2}\geq\left\lfloor(2K^{2}a_{k}a_{i}/a_{j}^{2})^{1/3}\right\rfloor$, which can be minorized by the same way. Hence, we have $R_{1}+R_{2}+R_{3}\geq R_{0}$ for $i=1,2,3$ for a constant $R_{0}$ in Table 2 in each case.
@@
(48) $C_{3}=\frac{g_{0}C_{0}}{\log\rho}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}\log^{2}B_{0}}+C_{3}^{\prime}\right)$
with the note that $C_{3}\leq C_{3,0}$ in Table 3.
Now we would like to show that $\log b\leq\log b_{0}$. We see that $R_{i}-1=R_{i1}+R_{i2}+R_{i3}<c_{i}^{\prime}a_{i+1}a_{i+2}\log b_{0}$ for $i=1,2,3$. Hence, we obtain
@@
proceeding as in the proof of Lemme 8 of [7] and observing that
(53) $\frac{2+6\pi^{-2}+\log K}{3K(K-1)}+\frac{1-2\log(2\pi K/e^{3/2})}{K-1}<0$
for $K\geq 2$. Since $\log b_{0}\geq\log B_{0}=20$ and $C_{1}>e^{3/2}/2$ by assumption, we have
@@
(59) $\left(1-\frac{1}{K}\right)\left(1-\frac{1}{L}-\frac{1}{3L^{2}}\right)C_{0}C_{1}\geq\frac{2\log(C_{0}(C_{1}a_{1}a_{2}a_{3})^{2}\log^{3}b_{0}/\log\rho)}{a_{1}a_{2}a_{3}\log^{2}b_{0}}+C_{1}+C_{3,0}$
@@
(77) $\begin{split}&\rho_{1}\left\lvert\log\alpha_{i}^{\prime}\right\rvert+2\operatorname*{h}(\alpha_{i}^{\prime})<u_{1}\left(\log m_{i}+\frac{\rho_{1}\pi}{2}\right)+\frac{u_{i+1}\rho_{1}\pi}{2}\\ &<(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}\left(\left(\log m_{i}+\frac{\rho_{1}\pi}{2}\right)\frac{\rho\pi}{2}+\left(\log m_{i}+\frac{\rho\pi}{2}\right)\frac{\rho_{1}\pi}{2}\right)\\ &=(1+\epsilon)\chi_{1}^{-2/3}C_{1}^{1/3}((\rho+\rho_{1})\log m_{i}+\rho\rho_{1}\pi)\pi/2.\end{split}$
@@
(92) $\begin{split}{b^{\prime}}=&\frac{b_{1}}{9.05\pi+2D\operatorname*{h}(\alpha)}+\frac{b_{2}}{9.05\pi},\\ D=&[{\mathbb{Q}}(\alpha):{\mathbb{Q}}]/2,\\ a=&9.05\pi+2D\operatorname*{h}(\alpha),\\ H=&\max\{17,D,D(\log{b^{\prime}}+2.96)+0.01\}.\end{split}$
@@
(93) $\log\left\lvert\Lambda_{1}\right\rvert>-2.7717aH^{2}.$
@@
Then, we can confirm that $0.27470.2432<\sqrt{k}<$ 0.2747, $H>7.5$, $15\leq L<0.92h$, and $K>700$ and then the condition (2) of Theorem 1 of [6]. Moreover, we can also confirm that $R_{2}-1<2a_{2}H^{2}<b_{1}$ and $S_{2}-1<2a_{1}H^{2}<b_{2}$. Hence the condition (1) of Theorem 1 of [6] also holds. Then, confirming that $\sqrt{k}L<0.23865H$, Theorem 1 of [6] yields that $\log\left\lvert\Lambda_{1}\max\{LSe^{LS\left\lvert\Lambda_{1}\right\rvert/(2b_{2})}/(2b_{2}),LRe^{LR\left\lvert\Lambda_{1}\right\rvert/(2b_{1})}/(2b_{1})\}\right\rvert>-2.7706aH^{2}$ and then $\log\left\lvert\Lambda_{1}\right\rvert>-2.7717aH^{2}$. For the detail of the proof, see [22]. ∎
@@ 5. Proof of the Theorem
If $m_{2}\geq 10^{34}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{34})$ again, we have $\left\lvert y_{i}\right\rvert<9.098\times 10^{16}$ and $\log x_{i}<8380114833$. If $10^{32}\leq m_{2}<10^{34}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{32})$, we have $\left\lvert y_{i}\right\rvert<9.152\times 10^{16}$ and $\log x_{i}<8373165769$. If $10^{24}\leq m_{2}<10^{32}$, then, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(53,10^{24})$, we have $\left\lvert y_{i}\right\rvert<8.752\times 10^{16}$ and $\log x_{i}<7706422610$. If $m_{2}<10^{24}$, then, applying Theorem 4.1 again with $\tilde{m}_{1}=53$ and $\tilde{m}_{2}=m_{1}^{*}$, we have $\left\lvert y_{i}\right\rvert<7.799\times 10^{16}$ and $\log x_{i}<5456058110$. Hence, in any case, we have $y_{i}<9.152\times 10^{16}$ and $x_{i}<\exp(8380114833)$.
@@
We begin by showing that $m_{1}<M_{1}$ with $M_{1}=2.549\times 10^{24}$. Assume that $m_{2}\geq m_{1}\geq M_{1}$. Now, Theorem 4.1 allows us to take $A_{4}=1752.029$, $A_{5}=10545512.8\cdots$, and $\rho=7.292$.
@@
Now, if $m_{1}\leq 6.5\times 10^{8}G^{2}$, then, (120) gives $m_{1}<6.5\times 10^{8}A_{4}^{2}(\log^{2}m_{1})\log^{4}(10545513\log m_{1})$ and therefore we must have $m_{1}<2.549\times 10^{24}<M_{1}$. Thus, we may assume that $m_{2}>m_{1}>6.5\times 10^{8}G^{2}$.
@@
(124) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{m_{1}}}>(1-10^{-9})\left(\frac{\pi}{4}-7.9\times 10^{-5}\right)>0.785319.$
@@
(129) $\log\left\lvert\Lambda_{1}\right\rvert>-2.7717aH^{2},$
@@
(130) $\begin{split}a=&9.05\pi+\log m_{1}\\ H=&\max\left\{17,2.97+\log\left(\frac{\left\lvert r\right\rvert}{9.05\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9.05\pi}\right)\right\}.\end{split}$
@@
Since $m_{1}\geq M_{1}$, we have $0.3774942.97+\log\psi_{3}(m_{1})\leq 2.97+\log\psi_{3}(M_{1})<-$ 0.377494. Hence, if $H>17$, then
@@
We observe that the right hand side of this inequality is at least
$-0.377494+\log(A_{4}\log M_{1})\log^{2}(10545513\log M_{1})>17$ to obtain
@@
whether $H>17$ or not. With the aid of these inequalities, we deduce from (128) and (129) that
@@
(139) $A_{4}(\alpha\log m_{2})\log^{2}(10545513\alpha\log m_{1})<0.785318\sqrt{m_{1}}$
@@
Now we show that $m_{2}<\exp(295490)$. Assume that $m_{2}\geq\exp(295490)$. If $m_{1}\geq M_{1}^{\prime}:=9.203\times 10^{22}$, then, since $m_{2}>e^{40000}$, we iterate the argument that we have just done with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ to obtain (137) and (138) with $42.29385$, $116.73985$, and $441.0489$ in place of $150.9763$, $381.3734$, and $1922.630$. Eventually, we obtain
(140) $A_{4}(\alpha\log m_{2})\log^{2}(6665783\alpha\log m_{1})<0.785358\sqrt{m_{1}}$
with $A_{4}=1675.249$ instead of (139). On the other hand, if $m_{1}\leq 5\times 10^{14}G^{2}$, then, (120) gives $G<A_{4}(\log^{2}(5\times 10^{14}G^{2}))\log^{4}(6665783\log(2.56\times 10^{10}G^{2}))$ and therefore $G<3\times 10^{12}$, which implies that $m_{2}>e^{40000}>5\times 10^{14}G^{2}$. Hence, in any case, we must have $m_{2}>5\times 10^{14}G^{2}$ to obtain (123) and then (126) with $10^{-7}$ in place of $7.9\times 10^{-5}$ and $0.785398$ in place of $0.785319$, contrary to (140). Hence, we must have $m_{1}<M_{1}^{\prime}$ when $m_{2}\geq e^{40000}$.
@@
Moreover, if $5\leq m_{1}\leq 41$ and $m_{2}$ is at least the constant in the corresponding row in Table 6, then we obtain (136) with $A_{4}$ taken from Table 4 with $\tilde{m}_{2}=e^{40000}$, which is incompatible with (126).
@@
Hence, we must have $m_{1}<2.549\times 10^{24}$ and $m_{2}<\exp(295490)$. Moreover, if $m_{1}\geq 9.203\times 10^{22}$, then $m_{2}<e^{40000}$. If $m_{1}\geq 10^{14}$ and $m_{2}\geq e^{40000}$, then, Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ yields that $\log x\leq(\log 2+k_{i}\log m_{1}+\ell_{i}\log m_{2})/2<5.07\times 10^{12}$, $x_{i}<\exp(5.07\times 10^{12})$, and $\left\lvert y_{i}\right\rvert\leq 2FG<1.316\times 10^{19}$. Similarly, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})$ according to Table 6 in each case, we can confirm that $x_{i}<\exp(5.07\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<1.319\times 10^{19}$ whenever $m_{1}\geq 53$.